OpenStudy (anonymous):

find the volume of a cone whose vertex is the center of a sphere of radius 5 and whose base is the intersection of this sphere with a plane 3 units away from the sphere's center

OpenStudy (anonymous):

so diameter of the base is sqrt(5^2-3^2)=4 so volume \[(1/3)\pi (4)^{2}*3=16\pi\]

OpenStudy (anonymous):

thanks